# DHP Benchmark: Are LLMs Good NLG Evaluators?

### v1

@@ A.1 Summarization
We utilize the SummEval Fabbri et al. (2021) (MIT license) and SumPubMed Gupta et al. (2020) datasets (MIT license) for our summarization tasks. The SummEval dataset comprises 100 news articles, each accompanied by multiple reference and generated summaries. For our analysis, we exclusively use the reference summaries, selecting the one with the highest number of sentences from each article to facilitate perturbation. The SumPubMed dataset contains 32,000 long scientific articles along with their abstracts serving as reference summaries. We only use the "BACKGROUND" sections of these articles and summaries. From this dataset, we randomly select 100 pairs of articles and their corresponding summaries.
@@ Appendix B Hierarchical Perturbation
The specifics of the hierarchical perturbations are detailed in Table 2. We perform these perturbations based on character, word, and sentence-level statistical data of the texts, which are also presented in Table 2. Our rule-based perturbations include simple text deletions, typographical errors using existing software tools, reordering of sentences, and the incorporation of random or incorrect sentences from other data.
@@ Appendix C Expert Weights
[…]

### v2

@@ A.1 Summarization
We utilize the SummEval Fabbri et al. (2021) (MIT license) and SumPubMed Gupta et al. (2020) datasets (MIT license) for our summarization tasks. The SummEval dataset comprises 100 news articles, each accompanied by multiple reference and generated summaries. For our analysis, we exclusively use the reference summaries, selecting the one with the highest number of sentences from each article to facilitate perturbation. The SumPubMed dataset contains 32,000 long scientific articles along with their abstracts serving as reference summaries. We only use the "BACKGROUND" sections of these articles and summaries. We randomly select 100 pairs of articles and their corresponding summaries.
@@ A.4 Translation
Table 2: Summary of hierarchical perturbation methods applied to different NLG tasks, detailing the types of perturbations and their respective implementations based on character (C), word (W), and sentence-level (S) modification with rule-based (R) or LLM-based (L) approaches.
Task Avg NLTK Statistics Perturbation Description
Summarization SummEval: 340.4 Characters 58.3 Words 4.0 Sentences SumPubMed 803.5 Characters 114.9 Words 5.5 Sentences (C, R) Random Deletions
Delete k alphanumeric characters randomly.
SummEval: k=10 for Minor, k=50 for Major;
SumPubMed: k=20 for Minor, k=100 for Major.
(C, R) Random Typos
Add k random typographical errors with "typo" package.
SummEval: k=10 for Minor, k=50 for Major;
SumPubMed: k=20 for Minor, k=100 for Major.
(W, L) Fictional Named Entities
Substitute one ore more named entities with in the
summary (e.g., names, locations, specific numbers,
technical terms, etc.) with fictional counterparts.
(W, L) Grammatical Errors
Modify the summary for creating two or more
grammatical errors, such as subject-verb
disagreement, noun-pronoun disagreement,
incorrect verb tense, misuse of preposition,
and sentence fragment, etc.
(S, R) Reordering
Random shuffle k sentences in the summary.
k=2 for Minor, k=all for Major.
(S, L) Rewriting and Insertion
Select one or more sentences from the summary,
then rephrase them and insert the rewritten
versions immediately after the original sentences.
Story Completion Story Cloze Test: 38.7 Characters 7.4 Words 1.0 Sentences (C, R) Random Deletions Delete 5 alphanumeric characters randomly.
(C, R) Random Typos
Add 5 random typographical errors with "typo" package.
(W, L) Fictional Named Entities
Substitute one critical named entities within the ending
sentence (e.g., a name, a location, a specific number, etc.)
with a fictional counterpart.
(W, L) Grammatical Errors
Modify the ending for creating one grammatical error,
such as subject-verb disagreement, noun-pronoun
disagreement, incorrect verb tense, misuse of preposition,
and sentence fragment, etc.
(S, R) Random Ending Sentence Replace the ending with a random one from another story.
(S, R) Wrong Ending Sentence Replace the ending with the wrong ending of the dataset.
Question Answering Answer Equivalence: 156.2 Characters 23.9 Words 1.0 Sentences (C, R) Random Deletions
Delete k alphanumeric characters randomly.
k=5 for Minor, k=25 for Major.
(C, R) Random Typos
Add k random typographical errors with "typo" package.
k=5 for Minor, k=25 for Major.
(W, L) Fictional Named Entities
Substitute one or more critical named entities within
the answer (e.g., names, locations, specific numbers,
technical terms, etc.) with fictional counterparts.
(W, L) Grammatical Errors
Modify the answer for creating one or more grammatical
errors, such as subject-verb disagreement, noun-pronoun
disagreement, incorrect verb tense, misuse of preposition,
and sentence fragment, etc.
(S, R) Random Answer Replace the answer with a random one to another question.
Translation WMT-22 German-to-English: 436.8 Characters 71.0 Words 3.8 Sentences WMT-22 Chinese-to-English: 434.1 Characters 66.4 Words 1.1 Sentences (C, R) Random Deletions
Delete k alphanumeric characters randomly.
k=10 for Minor, k=50 for Major.
(C, R) Random Typos
Add k random typographical errors with "typo" package.
k=10 for Minor, k=50 for Major.
(W, R) Random Deletions
Delete k continuous words in the translation randomly.
k=5 for Minor, k=25 for Major.
(W, L) Fictional Named Entities
Substitute one or more critical named entities within
the translation (e.g., names, locations, specific numbers,
technical terms, etc.) with fictional counterparts.
(W, L) Grammatical Errors
Modify the translation for creating two or more
grammatical errors, such as subject-verb disagreement,
noun-pronoun disagreement, incorrect verb tense, misuse
of preposition, and sentence fragment, etc.
@@ Appendix B Hierarchical Perturbation
The specifics of the hierarchical perturbations are detailed in Table 2. We perform these perturbations based on character, word, and sentence-level statistical data of the texts, which are presented in Table 2. Our rule-based perturbations include simple text deletions, typographical errors using existing software tools, reordering of sentences, and the incorporation of random or incorrect sentences from other data.
@@
Figure 5: User interface of the expert weight survey conducted to determine the impact of various quality issues on NLG task metrics.
@@ Appendix C Expert Weights
Model Version Source
GPT3.5-Turbo gpt-3.5-turbo-0125 platform.openai.com/docs/models